# KeepKV: Eliminating Output Perturbation in KV Cache Compression for Efficient LLMs Inference

### v1

@@ A.4 Error Bound Analysis
After extending KeepKV to multi-step generation in Section 3.3.2, for $t^{\prime}$-step, all $s_{i}^{t}$ terms in Equation 25 are replaced with $\hat{s_{i}^{t^{\prime}}}$ , which represents our estimation of future attention score trends obtained through a certain method. In this case, the merging expressions become:
$k_{r}=\frac{(\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}k_{i})\ln{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}\ln{\hat{s_{i}^{t^{\prime}}}}},\quad v_{r}=\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}v_{i}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}},\quad p_{r}=\sum_{i\in K^{\prime}_{e}}p_{i}$ (26)
@@
$\Theta_{t^{\prime}}$ $=\left\|o_{t^{\prime}}-o^{\prime}_{t^{\prime}}\right\|$ (28)
@@
$=\frac{\left\|\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(v_{j}-v_{i})-p_{r}s_{r}^{t^{\prime}}(v_{r}-v_{i})]\right\|}{(\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}})(\sum_{i=1,i\notin K^{\prime}_{e}}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}+p_{r}s_{r}^{t^{\prime}})}$
Substituting the expression for $v_{r}$ from Equation 26 and $\sum_{j\in K^{\prime}_{e}}p_{j}\hat{s_{j}^{t^{\prime}}}=p_{r}\hat{s_{r}^{t^{\prime}}}$ from Equation 27 into the above: s
$\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(v_{j}-v_{i})-p_{r}s_{r}^{t^{\prime}}(v_{r}-v_{i})]$ (29)
$=\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(v_{j}-v_{i})-p_{r}s_{r}^{t^{\prime}}(\frac{\sum_{k\in K^{\prime}_{e}}p_{k}\hat{s_{k}^{t^{\prime}}}v_{k}}{\sum_{k\in K^{\prime}_{e}}p_{k}\hat{s_{k}^{t^{\prime}}}}-v_{i})]$
$=\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}})(v_{j}-v_{i})]$
@@
$\left\|\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}})(v_{j}-v_{i})]\right\|$ $\leq\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}(\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}\left|1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}\right|\left\|v_{j}-v_{i}\right\|)$ (31)
$\leq\frac{2\epsilon\gamma}{1-\epsilon}(\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}})(\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}})$
@@
By Theorem A.2, it is easy to obtain that when $\epsilon=0$, $\Theta_{t^{\prime}}<\frac{2\epsilon(1+\epsilon)\gamma}{(1-\epsilon)^{2}}=0$. Next, we prove that when $(k_{i},v_{i})=(k_{j},v_{j}),\forall i,j\in K^{\prime}_{e}$, it also holds that $\Theta_{t^{\prime}}=0$. First, we further expand $(1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}),j\in K^{\prime}_{e}$ in Equation 29 by applying Equation 27:
@@
$1-\frac{e^{\frac{q_{t^{\prime}}k_{r}}{\sqrt{d}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}$ $=1-\frac{{(\prod_{i\in K^{\prime}_{e}}{s_{i}^{t^{\prime}}}^{p_{i}\hat{s_{i}^{t^{\prime}}}})^{\frac{\ln{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}\ln{\hat{s_{i}^{t^{\prime}}}}}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}$ (36)
@@
$1-\frac{{(\prod_{i\in K^{\prime}_{e}}{s_{i}^{t^{\prime}}}^{p_{i}\hat{s_{i}^{t^{\prime}}}})^{\frac{\ln{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}\ln{\hat{s_{i}^{t^{\prime}}}}}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s_{i}^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}$ $=1-\frac{{(\prod_{i\in K^{\prime}_{e}}{s^{t^{\prime}}}^{p_{i}\hat{s^{t^{\prime}}}})^{\frac{\ln{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s^{t^{\prime}}}\ln{\hat{s^{t^{\prime}}}}}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s^{t^{\prime}}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\frac{\hat{s^{t^{\prime}}}}{s^{t^{\prime}}}$ (37)
$=1-\frac{s^{t^{\prime}}}{\hat{s^{t^{\prime}}}}\frac{\hat{s^{t^{\prime}}}}{s^{t^{\prime}}}=0$
@@
$\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(v_{j}-v_{i})-p_{r}s_{r}^{t^{\prime}}(v_{r}-v_{i})]=\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}})(v_{j}-v_{i})]=0$ (38)
@@
This lemma provides a theoretical justification for prior merging strategies favoring high-similarity KV pairs [23, 24]. Meanwhile, we offer an intuitive interpretation: if the merged two KV pairs are identical, i.e., $(k_{e},v_{e})=(k_{c},v_{c})$, retaining one pair and setting its Electoral Votes to 2 introduces no error in subsequent computations.
@@
In this section, we have proven Theorem A.2 and Lemma A.3, which provide guarantees on the output perturbation in multi-step generation—an aspect that existing methods struggle to achieve. Moreover, our method demonstrates superior performance across various experimental evaluations. However, it should be acknowledged that predicting attention distributions further into the future is inherently challenging, leading to a significant increase in the estimated perturbation upper bound. Furthermore, the inherent input differences $\gamma$ cannot be ignored, representing a fundamental problem in KV cache compression—namely, the inability to perfectly compress the KV cache into a smaller memory without any loss of information. Nevertheless, our work introduces a new perspective and analytical approach to studying KV cache eviction and merging algorithms, which we hope will inspire future research.

### v2

@@ A.4 Error Bound Analysis
After extending KeepKV to multi-step generation, for $t^{\prime}$-step, all $s_{i}^{t}$ terms in Equation 25 are replaced with $\hat{s_{i}^{t^{\prime}}}$ , which represents our estimation of future attention score trends obtained through a certain method. In this case, the merging expressions become:
$k_{r}$ $=\frac{\left(\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}k_{i}\right)\ln{\left(\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}\right)}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}\ln{\hat{s}_{i}^{t^{\prime}}}},$
$v_{r}$ $=\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}v_{i}}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}},\quad p_{r}=\sum_{i\in K^{\prime}_{e}}p_{i}.$ (26)
@@
$\Theta_{t^{\prime}}=\left\|o_{t^{\prime}}-o^{\prime}_{t^{\prime}}\right\|$
@@
$=\frac{\left\|\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}\left(\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(v_{j}-v_{i})-p_{r}s_{r}^{t^{\prime}}(v_{r}-v_{i})\right)\right\|}{\left(\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}\right)\left(\sum_{i=1,\,i\notin K^{\prime}_{e}}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}+p_{r}s_{r}^{t^{\prime}}\right)}$ (28)
Substituting the expression for $v_{r}$ from Equation 26 and $\sum_{j\in K^{\prime}_{e}}p_{j}\hat{s_{j}^{t^{\prime}}}=p_{r}\hat{s_{r}^{t^{\prime}}}$ from Equation 27 into the above:
$\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}\left[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(v_{j}-v_{i})-p_{r}s_{r}^{t^{\prime}}\left(\frac{\sum_{k\in K^{\prime}_{e}}p_{k}\hat{s}_{k}^{t^{\prime}}v_{k}}{\sum_{k\in K^{\prime}_{e}}p_{k}\hat{s}_{k}^{t^{\prime}}}-v_{i}\right)\right]$
$=\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}\left(1-\frac{s_{r}^{t^{\prime}}}{\hat{s}_{r}^{t^{\prime}}}\cdot\frac{\hat{s}_{j}^{t^{\prime}}}{s_{j}^{t^{\prime}}}\right)(v_{j}-v_{i})$ (29)
@@
$\left\|\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}\left[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}\left(1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\cdot\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}\right)(v_{j}-v_{i})\right]\right\|$
$\leq\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}\left(\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}\left|1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\cdot\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}\right|\cdot\|v_{j}-v_{i}\|\right)$
$\leq\frac{2\epsilon\gamma}{1-\epsilon}\left(\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}\right)\left(\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}\right)$ (31)
@@
By Theorem 8, it is easy to obtain that when $\epsilon=0$, $\Theta_{t^{\prime}}<\frac{2\epsilon(1+\epsilon)\gamma}{(1-\epsilon)^{2}}=0$. Next, we prove that when $(k_{i},v_{i})=(k_{j},v_{j}),\forall i,j\in K^{\prime}_{e}$, it also holds that $\Theta_{t^{\prime}}=0$. First, we further expand $(1-\frac{s_{r}^{t^{\prime}}}{\hat{s_{r}^{t^{\prime}}}}\frac{\hat{s_{j}^{t^{\prime}}}}{s_{j}^{t^{\prime}}}),j\in K^{\prime}_{e}$ in Equation 29 by applying Equation 27:
@@
$1-\frac{e^{\frac{q_{t^{\prime}}k_{r}}{\sqrt{d}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\cdot\frac{\hat{s}_{j}^{t^{\prime}}}{s_{j}^{t^{\prime}}}$
$=1-\frac{\left(\prod_{i\in K^{\prime}_{e}}(s_{i}^{t^{\prime}})^{p_{i}\hat{s}_{i}^{t^{\prime}}}\right)^{\frac{\ln\left(\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}\right)}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}\ln\hat{s}_{i}^{t^{\prime}}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\cdot\frac{\hat{s}_{j}^{t^{\prime}}}{s_{j}^{t^{\prime}}}$ (36)
@@
$1-\frac{\left(\prod_{i\in K^{\prime}_{e}}{s_{i}^{t^{\prime}}}^{p_{i}\hat{s}_{i}^{t^{\prime}}}\right)^{\frac{\ln\left(\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}\right)}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}\ln\hat{s}_{i}^{t^{\prime}}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}_{i}^{t^{\prime}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\cdot\frac{\hat{s}_{j}^{t^{\prime}}}{s_{j}^{t^{\prime}}}$
$=1-\frac{\left(\prod_{i\in K^{\prime}_{e}}{s^{t^{\prime}}}^{p_{i}\hat{s}^{t^{\prime}}}\right)^{\frac{\ln\left(\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}^{t^{\prime}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}\right)}{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}^{t^{\prime}}\ln\hat{s}^{t^{\prime}}}}}{\frac{\sum_{i\in K^{\prime}_{e}}p_{i}\hat{s}^{t^{\prime}}}{\sum_{i\in K^{\prime}_{e}}p_{i}}}\cdot\frac{\hat{s}^{t^{\prime}}}{s^{t^{\prime}}}$
$=1-\frac{s^{t^{\prime}}}{\hat{s}^{t^{\prime}}}\cdot\frac{\hat{s}^{t^{\prime}}}{s^{t^{\prime}}}=0$ (37)
@@
$\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}\left[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}(v_{j}-v_{i})-p_{r}s_{r}^{t^{\prime}}(v_{r}-v_{i})\right]$
$=\sum_{i=1}^{t^{\prime}}p_{i}s_{i}^{t^{\prime}}\left[\sum_{j\in K^{\prime}_{e}}p_{j}s_{j}^{t^{\prime}}\left(1-\frac{s_{r}^{t^{\prime}}}{\hat{s}_{r}^{t^{\prime}}}\cdot\frac{\hat{s}_{j}^{t^{\prime}}}{s_{j}^{t^{\prime}}}\right)(v_{j}-v_{i})\right]=0$ (38)
@@
This lemma provides a theoretical justification for prior merging strategies favoring high-similarity KV pairs (wan2024D2O; wang2024KVMerger). Meanwhile, we offer an intuitive interpretation: if the merged two KV pairs are identical, i.e., $(k_{e},v_{e})=(k_{c},v_{c})$, retaining one pair and setting its Electoral Votes to 2 introduces no error in subsequent computations.
@@
In this section, we have proven Theorem 8 and Lemma 9, which provide guarantees on the output perturbation in multi-step generation—an aspect that existing methods struggle to achieve. Moreover, our method demonstrates superior performance across various experimental evaluations. However, it should be acknowledged that predicting attention distributions further into the future is inherently challenging, leading to a significant increase in the estimated perturbation upper bound. Furthermore, the inherent input differences $\gamma$ cannot be ignored, representing a fundamental problem in KV cache compression—namely, the inability to perfectly compress the KV cache into a smaller memory without any loss of information. Nevertheless, our work introduces a new perspective and analytical approach to studying KV cache eviction and merging algorithms, which we hope will inspire future research.
Appendix B Implementation Details
B.1 Models and benchmarks
Across all experiments, we utilized pre-trained model weights from Huggingface (wolf2020huggingfacestransformers). Specifically, for Llama architectures: we used the ’huggyllama/llama-7b’ checkpoint for Llama-1-7B (touvron2023llama), ’meta-llama/Llama-2-7b-hf’ for Llama-2-7B (touvron2023llama2), and ’meta-llama/Llama-2-13b-hf’ for Llama-2-13B. For Llama-3-8B, the ’meta-llama/Meta-Llama-3-8B-Instruct’ checkpoint was used in the HELM evaluations, and ’meta-llama/Meta-Llama-3-8B’ in LongBench evaluations. Regarding the Mistral architecture, we employed the ’mistralai/Mistral-7B-Instruct-v0.3’ model (jiang2023mistral7b). For detailed information on the LongBench benchmark, please refer to its official repository (bai2024longbench).
B.2 Experimental Setup
Our study does not involve model training, thus no data preprocessing is required. All evaluation datasets are sourced from the publicly available lm-eval-harness (lmevalharness), HELM (narayan2018xsum; nallapati2016cnndm) and LongBench (bai2024longbench) benchmarks, and we follow their original evaluation metrics. Our implementation is primarily based on modifications to the open-source H2O and D2O codebases (zhang2023h2o; wan2024D2O), both of which are publicly accessible. All open-source datasets, evaluation frameworks, and algorithmic implementations used in this work are employed in full compliance with their respective licenses and terms of use.
B.3 Details of Parameter Settings
For cache allocation, we adopt the default configuration provided by the open-source PyramidInfer (yang2024pyramidinfer) implementation. Specifically, we follow a cosine-based decay strategy for distributing cache across layers. Within each layer, the ratio between the recent window and the heavy hitter (i.e., crucial KV entries) varies between 5:1 and 4:1. Additionally, we retain the first 4 tokens (referred to as ’sink tokens’ in StreamingLLM (xiao2024StreamingLLM)) in the cache throughout. For algorithmic hyperparameters, we adjust the merging threshold $T$ within the range of 0.6 to 0.95, and empirically select 0.8 as the default value based on its stable performance across benchmarks. The corresponding experimental results are reported below.
Prefill’s Merge Policy
Cache Budget 1% 10%
Evict-first 0.051 0.111
Merge-first 0.059 0.115
Table 3: Performance comparison with different merge policy during the prefill stage (ROUGE-2; LLAMA-7B).
The results in Table 3 show that, during the prefill stage, performing merging before applying the eviction strategy—rather than merging after eviction as is commonly done—can improve generation quality. This is because, according to Lemma 6, merging more similar items induces smaller perturbations. In the prefill stage, there exist more tokens whose similarity exceeds the threshold; merging them first allows these tokens to be consolidated rather than evicted unnecessarily, thereby preserving more information within the limited-size cache.
Merging Threshold
Merging Thres XSUM
0.7 0.223
0.8 0.233
0.9 0.214
Table 4: Merging threshold impact (4% compression ratio).
We investigate the impact of different merging thresholds on performance. The results in Table 4 indicate that setting the merging threshold (i.e., the cosine similarity between key vectors) to 0.8 yields better performance, and its robustness has been further validated across a broader range of experiments.
Appendix C Additional Discussion
C.1 Limitations and Future Work
As analyzed in Section A, the bound on output perturbation in multi-step generation depends on two quantities: the intrinsic difference of the inputs $\gamma$, and the prediction error $\epsilon$. The former is inherent and cannot be influenced or altered, whereas the latter is inevitably non-negligible in realistic inference scenarios. In practice, accurately predicting future attention scores is highly challenging. Prior work typically employs sliding-window averages or weighted averages as predictors, while more recent studies have begun to explore machine-learning-based approaches to obtain more accurate predictions; this appears to be a promising direction for future research.
Moreover, the merging operation inevitably incurs additional computational overhead, which can render it less favorable than purely eviction-based strategies with simpler computation in certain settings. Our mathematically derived lossless KV cache merging method, along with many other related algorithms, further relies on access to attention scores, which limits its compatibility with widely adopted inference acceleration techniques such as FlashAttention. Future work could therefore focus on designing KV cache eviction and merging algorithms that are amenable to integration with FlashAttention and related acceleration methods.
C.2 Comparison with CaM in Eliminating Output Perturbations
A closely related line of work is CaM (zhang2024cam), which also analyzes the effect of KV cache eviction and its proposed merging algorithm on output perturbation, and aims to reduce or even eliminate such perturbation. However, there are two fundamental differences between CaM and our approach: (i) how ”output perturbation” is defined and modeled, and (ii) how the theoretical analysis is connected to the asynchronous KV cache updates that actually occur in autoregressive decoding.
Before presenting our analysis, it is important to clarify the implementation setting shared by both CaM and our method. In practical autoregressive decoding, KV compression is always performed inside the decoding loop: the model first runs the forward pass at step $t$ using the current KV cache and produces the output, and only afterwards applies eviction or merging operations to the cache state that will be consumed at step $t+1$ and beyond. In other words, cache compression is asynchronous with respect to the forward computation. Consequently, when we design an algorithm from the perspective of eliminating output perturbation, our first requirement is consistency at the current step: if we hypothetically recompute the step-$t$ forward pass on the compressed KV cache, the resulting output should match that obtained from the original, uncompressed cache. Intuitively, the attention computation between the step-$t$ query and the pre-compression KV cache should yield exactly the same result as the attention between the same query and the post-compression KV cache. This ensures that, when we ”look back” at step $t$ in the future, its behavior is indistinguishable from a world in which step $t$ had always been computed on the compressed KV cache, without needing to refer to the original KV entries that are no longer stored. The overall decoding process thus appears temporally coherent.
In contrast, when CaM studies the effect of eviction and merging on the output at step $t$, it does not adopt such a ”recompute on the compressed cache” viewpoint. Instead, it directly removes (or rescales) the contribution of the evicted or merged KV pair in the attention scores and in the attention-weighted output, while leaving the contributions of all remaining tokens unchanged. This effectively treats the attention scores of the other KV pairs as fixed coefficients that are not affected by modifying the cache. However, due to the softmax normalization, evicting or merging any KV pair inevitably changes the normalization term and thus the attention weights of all remaining KV pairs. Capturing this global effect of softmax is precisely why our analysis starts from the full attention expression, rather than from a model in which the other attention weights are held constant. In summary, CaM and our method formalize and analyze output perturbation from different perspectives, and each seeks to reduce or eliminate perturbation under its own definition, but our framework is explicitly aligned with the asynchronous KV-update pattern and with the impact of softmax on all tokens in the cache.